# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

### v2

@@ Abstract
Sparse phase retrieval, which aims to recover a $k$-sparse signal from $m$ phaseless measurements, poses a fundamental question regarding the minimal sample complexity required for success. While the optimal sample complexity is known to be $\Omega(k\log n)$, existing algorithms can only achieve it for signals under restrictive structural assumptions. This paper introduces a novel and robust initialization algorithm, termed generalized Exponential Spectral Pursuit (gESP), designed to overcome this limitation. Theoretically, we prove that gESP significantly expands the family of signals that can be recovered with the optimal sample complexity. Our analysis unifies existing results on previously studied signal models and surpasses them by establishing performance bounds for a more general class of signals. Extensive simulations validate our theoretical findings, demonstrating that gESP consistently outperforms state-of-the-art methods across diverse signal types, thereby pushing the boundaries of efficient and optimal sparse phase retrieval.
Index Terms:
Sparse phase retrieval, sample complexity, sparsity, information-theoretical bound.
@@ I Introduction
In many applications, only the intensity information is retained by the measurement system due to hardware limitations [1]. For example, in optical imaging [2, 3] and X-ray crystallography [4, 5], common sensors such as charge-coupled devices, photosensitive films and human eyes merely record the amplitude of the light wave but ignore the phase. In these settings, one is faced with the task of recovering the original signal from its magnitude-only measurements, which is often referred to as phase retrieval (PR) [6, 7, 8, 9, 10]. Succinctly stated, the main task of PR is to reconstruct the $n$-dimensional signal $\mathbf{x}\in\mathbb{C}^{n}$ from a group of phaseless measurements $\mathbf{y}\in\mathbb{R}^{m}$ with
@@
where $\mathbf{a}_{i}\in\mathbb{C}^{n}$ is the measurement vector. In this paper, we consider the complex Gaussian measurements, i.e., $\mathbf{a}_{i}\sim\mathcal{CN}(n)$, and $\mathbf{x}$ is $k$-sparse, which means that there are at most $k$ non-zero entries in $\mathbf{x}$. This problem is the so-called sparse phase retrieval problem. Researchers are interested in the minimum measurement number that one needs to reconstruct the signal [11, 12, 13, 14]. It has been shown that with m=4k2m=4k-2 phaseless measurements [15], the non-linear system (1) with a $k$-sparse prior produces the unique solution based on generic sensing measurements. Moreover, researches have proved that
@@
For polynomial algorithms, they appear to require more measurements for the exact recovery. Back in 2013, Li and Voroninski [18] utilized the matrix lifting technique and convex relaxation methods to solve the sparse phase retrieval problem. They proved that it requires
@@
measurements to produce an exact reconstruction. Moreover, they further proved that this requirement cannot be reduced to be smaller than $m=\Omega({k^{2}}/{\log^{2}n})$ for their supposed model. While their sample complexity revealed a clear discrepancy compared to (2), the authors were unsure whether this difference was inherent to the problem itself or simply a consequence of limitations in their modeling approach. This is afterwards called the statistical-to-computational gap.
This gap is further investigated in the thresholded Wirtinger Flow (TWF) algorithm [19], which is a typical greedy algorithm containing two stages — the initialization stage and the refinement stage. They showed that both stages require $m=\Omega(k^{2}\log n)$ samples for the exact final recovery. This two-stage approach was followed in subsequent algorithms, e.g., SPARse Truncated Amplitude flow (SPARTA[20], Compressive Phase Retrieval with Alternating Minimization (CoPRAM[21], and Subspace Phase Retrieval (SPR[22]. It is shown via some geometric analysis [23, 22, 24] and sophisticated optimization approaches [25] that the sample complexity of the refinement stage can be improved to the information-theoretical bound (2), if a $\delta$-neighborhood estimate of the target signal is provided. However, to achieve this, the initialization stage requires
@@
Despite the persistence of this gap, numerous efforts are underway to achieve improved results. For the sake of notational simplicity, we define $s(p)$ as
@@
where $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. A formal definition and the corresponding interpretation are presented in Section III-A. Wu and Rebeschini [27] presented the algorithm call Hadamard Wirtingle Flow (HWF). They showed that the gap vanished for some specific structure of the target signal $\mathbf{x}$. Specifically, by assuming that the minimum magnitude of non-zeros $x_{(k)}$ is on the order of $\frac{\|\mathbf{x}\|}{\sqrt{k}}$, i.e.,
@@
they show that the sample complexity can be reduced to
@@
This result (7) is further improved by Cai et al. [26] and Xu et al. [28]. In particular, Cai et al. [26] proposed Truncated Power method (TPM) to remove the requirements on the minimum magnitude of non-zeros (i.e., (6)). Xu et al. [28] introduce an exponential spectrum and proposed Exponential Spectral Pursuit (ESP) algorithm to eliminate the $\log^{3}n$ term in (7). Finally, the sample complexity is reduced to
$m=\Omega\left(ks(1)\log n\right)$ (8)
samples are sufficient for ESP to produce a good estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ without any assumption on $x_{(k)}$. However, this result only achieves the optimal bound (2) for a restrictive class of signals, i.e., those with a single dominant entry where $s(1)=\Theta(1)$. This limitation motivates the development of algorithms that can achieve optimal sample complexity for a broader class of signals.
In this paper, we develop the generalized Exponential Spectral Pursuit (gESP) algorithm to address this challenge. Specifically, gESP selects an index set rather than a single index in its first step (i.e., Step 1 in Table 1). This key modification allows gESP to achieve the optimal sample complexity for a much broader class of signals. Specifically, it relaxes the condition for optimality from the restrictive $s(1)=\Theta(1)$ to the more general $s(\lceil\sqrt{k}\rceil)=\Theta(1)$. We show that when the signal structure $s(p)$ is known, gESP produces a $\delta$-neighborhood estimate with
@@
It can be shown that in the case where $s(1)=\Theta(\sqrt{k})$ and $s(\sqrt{k})=\Theta(1)$, this result will also be better than (8). The detailed discussion can be found in Section IV.
The remainder of this paper is organized as follows. In Section II, we review the development of the initialization stage and introduce our algorithm. Detailed interpretations are provided step by step. Section III contains theoretical results for the sample complexity of gESP in different cases and makes simple comparisons between these results. In Section IV, discussions are directed in several aspects to illustrate the superiority of our result. In Section V, the proofs of the main theorems are given based on the analysis for each step of gESP. Numerical simulations and analysis are conducted in Section VI. Finally, we conclude our paper in Section VII.
@@ II Algorithms
We introduce some useful notations. Throughout the paper, we use lowercase and uppercase boldface letters to represent vectors and matrices, while employing normal font for real numbers. Let $[n]$ denote the set $\{1,2,3,\cdots n\}$. We denote $x_{(i)}$ as the $i$-largest entry of $\mathbf{x}$ in magnitude. For the sake of readability, we denote $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ as the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. For any set $\subset[n], S{\mathbb{C}}^{S}$ is defined as the subspace of ${\mathbb{C}}^{n}$ spanned by vectors supported on $S$, i.e., $\{\mathbf{x}|\mathbf{x}\in{\mathbb{C}}^{n},{\text{supp}(\mathbf{x})}=S\}$. Unless otherwise specified, we define $\mathbf{a}_{S}$ as the vector that keeps the entries of $\mathbf{a}$ indexed by $S$ while setting others to zero. For the matrix $\mathbf{A}$, we define $\mathbf{A}_{S}$ as the matrix which keeps columns and rows indexed by $S$ while setting others to zero. Given a vector $\mathbf{x}$, the conjugate transpose, $\ell_{2}$ norm and $\ell_{0}$ norm of $\mathbf{x}$ are denoted as $\mathbf{x}^{*}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{0}$ respectively. An $n$-dimensional standard complex Gaussian random vector, denoted as $\mathbf{a}\sim\mathcal{CN}(n)$, is defined as $\mathbf{a}=\mathbf{a}_{1}+\mathrm{j}\mathbf{a}_{2}$, with $\mathbf{a}_{1},\mathbf{a}_{2}\sim\mathcal{N}(\mathbf{0},\frac{1}{2}\mathbf{I})$, and j is the imaginary unit. We use $a\sim b$ to represent that $a$ differs from $b$ by a constant factor. Following the convention in the literature, we use standard asymptotic notations to characterize the complexity as bellow.
@@
This assumption is made for the convenience of our brief introduction. One can find from our theorems that this assumption is actually unnecessary; also see the results in [26, 28].
@@ II-A Prior work
    Earlier researches contain two steps for initialization.
@@
    Generally, the expectation of $\mathbf{Z}$ takes the form as follow
@@
    where $\alpha,\beta$ are both numerical constants, determined by the specific definition of $Z_{j}$. For instance, [20, 29] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}|\mathbf{a}_{ij}|^{2}$ and the corresponding $\alpha=1,\beta=2$, whereas [14] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}|\mathbf{a}_{ij}|$ and the corresponding $\alpha=\frac{2}{\pi\|\mathbf{x}\|},\beta\approx\frac{1}{6\|\mathbf{x}\|}$. It can be seen that there exists an apparent gap between $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})}$ and $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})^{c}}$. Specifically,
    $\min_{j\in\operatorname*{supp}(\mathbf{x})}\mathbb{E}[Z_{j}]-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}\mathbb{E}[Z_{j}]=$ $\beta|{x_{(k)}}|^{2}$
    $\sim$ $\frac{\|\mathbf{x}\|^{2}}{k}.$ (14)
@@
    which means one can obtain the estimate of support set by sorting the elements of $\{Z_{j}\}_{j=1}^{n}$ and selecting the indices corresponding to the $k$ largest values. Then it can be guaranteed that this support estimate contains all the true candidates. Theoretical analysis shows that deriving an appropriate concentration from the gap $\frac{\|\mathbf{x}\|^{2}}{k}$ (see (14)) requires the sample complexity being [^1] When $Z_{j}$ is statistically heavy-tailed ((e.g., involving the forth power of Gaussian), some extra term containing $\log^{3}n$ (e.g., see (7)) will appear in the complexity. This is because one needs to use a truncation argument to analyze the tail bound. For simplicity, we will ignore such term throughout this subsection.
@@
    Step 2: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    The maximal eigenvector of $\mathbb{E}[\mathbf{Z}]$ is $\mathbf{x}$ and $\mathbb{E}[\|\mathbf{z}\|^{2}]=\|\mathbf{x}\|^{2}$. Therefore, we shall obtain a good estimate of $\mathbf{x}$. Theoretical analysis shows that this step requires
@@
    Recently, a novel initialization paradigm is proposed and developed in [27, 26, 28], which contains four steps. Instead of directly estimating the support in one step, an index $j_{\text{max}}$ will be firstly selected, which corresponds to the largest value among $Z_{j}$’s.
@@
    Establishing some concentration inequalities on this random variable yields the similar result to (15)
@@
    To identify the true support set for (21), it requires
@@
    Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    This step is as the same as Step 2 of the previous method.
@@ II-B Generalized Exponential Spectral Pursuit
As has been clarified, the enlarged gap $\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}$ leads to the improvement in complexity. Intuitively, the larger the gap is, the smaller the sample complexity will be. Note that the term $|{x_{(1)}}|$ in this gap arises from the fact that we simply select one maximal index in Step 1. If more significant indices are chosen in Step 1, the final gap will indeed be enlarged, and thus decrease the complexity in Step 3333In fact, the complexity in Step 1 will increase while that in Step 3 will decrease as more significant indices are chosen. This means that there is a trade-off in the number of the indices that should be chosen in Step 1.. Then the remaining issue is to design Step 2. Assume that we obtain the set $S^{0}$ from Step 1. Following the previous method, a straightforward idea is to construct the vector $\mathbf{e}^{0}$ such that it takes $1$ on $S^{0}$ and 0 for the rest. A more sophisticated design, however, is to compute the maximal eigenvector of $\mathbf{Z}_{S^{0}}$ as $\mathbf{e}^{0}$. This also matches the previous method since $\mathbf{e}_{j_{\text{max}}}$ is exactly the maximal eigenvector of $\mathbf{Z}_{\{j_{\text{max}}\}}$.
So far, we have completed the introduction of the modification idea. The specific algorithm is summarized in Algorithm 1. Now we will further clarify the effectiveness intuitively. The more detailed sample complexity and numerical conclusions are formally presented in Section III.
The spectrum in our proposed gESP algorithm is
@@
Since the maximal eigenvector of $\mathbb{E}[\mathbf{Z}_{S^{0}}]$ is $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$, we can approximate $\mathbf{e}_{0}$ as $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$. Then, Step 2 ensures that
@@
Therefore, Step 3 along with Step 4 ensure that $\mathbf{z}$ is a good initialization of the target signal $\mathbf{x}$.
Algorithm 1  Generalized Exponential Matching Pursuit
Input: $\mathbf{A}$, $\mathbf{y}$, the sparsity $k$, the parameter pp.
@@
Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
@@
Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.

### v3

@@ Abstract
Sparse phase retrieval aims to recover a $k$-sparse signal from $m$ phaseless measurements. While the theoretically optimal sample complexity for successful recovery is $\Omega(k\log n)$, existing algorithms can only achieve this bound for signals with specific structural assumptions, leading to a notable gap between theory and practice. To bridge this gap, we introduce an efficient initialization algorithm, termed generalized Exponential Spectral Pursuit (gESP). We prove that gESP can significantly expand the family of signals that are guaranteed to be recovered with the optimal sample complexity, thereby extending the scope of theoretical optimality to a much broader class of signals. Extensive simulations validate our theoretical findings and demonstrate that gESP consistently outperforms the state-of-the-art methods across diverse signal types.
@@ I Introduction
In many applications, only the intensity information is retained by the measurement system due to hardware limitations [2]. For example, in optical imaging [19, 18] and X-ray crystallography [16, 9], common sensors such as charge-coupled devices, photosensitive films, and human eyes merely record the amplitude of the light wave but ignore the phase. In these settings, one is faced with the task of recovering the original signal from its magnitude-only measurements, which is often referred to as phase retrieval (PR) [5, 35, 6, 24, 25]. Succinctly stated, the main task of PR is to reconstruct the $n$-dimensional signal $\mathbf{x}\in\mathbb{C}^{n}$ from a group of phaseless measurements $\mathbf{y}\in\mathbb{R}^{m}$ given by
@@
where $\mathbf{a}_{i}\in\mathbb{C}^{n}$ is the measurement vector. In this paper, we consider complex Gaussian measurements, i.e., $\mathbf{a}_{i}\sim\mathcal{CN}(n)$, and $\mathbf{x}$ is $k$-sparse, implying that there are at most $k$ non-zero entries in $\mathbf{x}$. This problem is known as the sparse phase retrieval problem. Researchers are interested in the minimum number of measurements required to reconstruct the signal [30, 31, 10, 17]. It has been shown that with m=4k2m=4k-2 phaseless measurements [28], the non-linear system (1) with a $k$-sparse prior produces a unique solution based on generic sensing measurements. Moreover, it has been proven that
@@
For polynomial-time algorithms performing exact recovery of $\mathbf{x}$, it appears to require significantly more phaseless measurements. Back in 2013, Li and Voroninski [14] utilized a matrix lifting technique and convex relaxation methods to solve the sparse phase retrieval problem. They proved that it requires
@@
measurements to produce an exact reconstruction. Moreover, they further proved that this requirement cannot be reduced to be smaller than $m=\Omega({k^{2}}/{\log^{2}n})$ for their supposed model. While their sample complexity revealed a clear discrepancy compared to (2), the authors were unsure whether this difference was inherent to the problem itself or simply a consequence of limitations in their modeling approach. Later, this discrepancy was identified as the statistical-to-computational gap [22, 21], particularly in the context of generic Gaussian measurements. Nevertheless, it has been shown that this barrier can be circumvented by imposing specific structures on the measurement process. Prior works, such as [1, 11], demonstrated that the optimal sample complexity is achievable in polynomial time when additional constraints are placed on the sensing vectors. However, under the standard unstructured Gaussian setting, verifying the existence of this gap remains a significant challenge.
This gap was further investigated in the thresholded Wirtinger Flow (TWF) algorithm [4], which is a typical greedy algorithm containing two stages: the initialization stage and the refinement stage. The authors showed that both stages require $m=\Omega(k^{2}\log n)$ samples for exact final recovery. This two-stage approach was followed in subsequent algorithms, e.g., SPARse Truncated Amplitude flow (SPARTA) [27], Compressive Phase Retrieval with Alternating Minimization (CoPRAM) [12], and Subspace Phase Retrieval (SPR) [32]. It is shown via some geometric analysis [21, 32, 15] and sophisticated optimization approaches [20] that the sample complexity of the refinement stage can be improved to the information-theoretical bound (2), provided that a $\delta$-neighborhood estimate of the target signal is available. However, to achieve this, the initialization stage requires
@@
Despite the persistence of this gap, numerous efforts are underway to achieve improved results. For the sake of notational simplicity, we define the structural function $s(p)$ as
@@
where $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ are the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes. A formal definition and the corresponding interpretation are presented in Section III-A. Wu and Rebeschini [29] presented an algorithm called Hadamard Wirtinger Flow (HWF). They showed that the gap vanished for signals with specific structures. Specifically, by assuming that the minimum magnitude of non-zeros $x_{(k)}$ is on the order of $\frac{\|\mathbf{x}\|}{\sqrt{k}}$, i.e.,
@@
they showed that the sample complexity can be reduced to
@@
The result of (7) was further improved by Cai et al. [3] and Xu et al. [33]. In particular, Cai et al. [3] proposed the Truncated Power Method (TPM) to remove the requirements on the minimum magnitude of non-zeros (i.e., Eq. (6)). Xu et al.  [33] introduced an exponential spectrum and proposed the Exponential Spectral Pursuit (ESP) algorithm to eliminate the $\log^{3}n$ term in (7). The sample complexity was reduced to
$m=\Omega\left(ks(1)\log n\right),$ (8)
which is sufficient for ESP to produce a good estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ without any assumption on $x_{(k)}$. However, this result only achieves the optimal bound (2) for a restrictive class of signals, i.e., those with a single dominant entry where $s(1)=\Theta(1)$. This limitation motivates the development of algorithms that can achieve optimal sample complexity for a broader class of signals.
In this paper, we develop the generalized Exponential Spectral Pursuit (gESP) algorithm to address this challenge. Specifically, gESP selects an index set rather than a single index in its first step (i.e., Step 1 in Algorithm 1). This key modification allows gESP to achieve the optimal sample complexity for a much broader class of signals. Specifically, it relaxes the condition for optimality from the restrictive $s(1)=\Theta(1)$ to the more general $s(\lceil\sqrt{k}\rceil)=\Theta(1)$. We show that when the signal structure $s(p)$ is known, gESP produces a $\delta$-neighborhood estimate with
@@
It can be shown that in the case where $s(1)=\Theta(\sqrt{k})$ and $s(\sqrt{k})=\Theta(1)$, this result will also be better than (8). A detailed discussion can be found in Section IV.
The remainder of this paper is organized as follows. In Section II, we review the development of the initialization stage and introduce our algorithm. Detailed interpretations are provided step by step. Section III contains theoretical results for the sample complexity of gESP in different cases and makes simple comparisons between these results. In Section IV, we discuss various aspects to illustrate the superiority of our result. In Section V, the proofs of the main theorems are given based on the analysis for each step of gESP. Numerical simulations and analysis are conducted in Section VI. Finally, we conclude our paper in Section VII.
@@ II Algorithms
We introduce some useful notations. Throughout the paper, we use lowercase and uppercase boldface letters to represent vectors and matrices, while employing normal font for real numbers. Let $[n]$ denote the set $\{1,2,3,\cdots n\}$. For the sake of readability, we denote $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ as the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes. For any set $\subset[n], S{\mathbb{C}}^{S}$ is defined as the subspace of ${\mathbb{C}}^{n}$ spanned by vectors supported on $S$, i.e., $\{\mathbf{x}|\mathbf{x}\in{\mathbb{C}}^{n},{\text{supp}(\mathbf{x})}=S\}$. Unless otherwise specified, we define $\mathbf{a}_{S}$ as the vector that keeps the entries of $\mathbf{a}$ indexed by $S$ while setting others to zero. For the matrix $\mathbf{A}$, we define $\mathbf{A}_{S}$ as the matrix which keeps columns and rows indexed by $S$ while setting others to zero. Given a vector $\mathbf{x}$, the conjugate transpose, $\ell_{2}$ norm, and $\ell_{0}$ norm of $\mathbf{x}$ are denoted as $\mathbf{x}^{*}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{0}$ respectively. An $n$-dimensional standard complex Gaussian random vector, denoted as $\mathbf{a}\sim\mathcal{CN}(n)$, is defined as $\mathbf{a}=\mathbf{a}_{1}+\mathrm{j}\mathbf{a}_{2}$, with $\mathbf{a}_{1},\mathbf{a}_{2}\sim\mathcal{N}(\mathbf{0},\frac{1}{2}\mathbf{I})$, and j is the imaginary unit. We use $a\sim b$ to represent that $a$ differs from $b$ by a constant factor. Throughout this paper, $c$ and $C$ denote universal positive constants whose values may change from line to line. We say an event occurs with high probability if it holds with probability at least $1-n^{-c}$ for some constant $c>0$. Following the convention in the literature, we use standard asymptotic notations to characterize the complexity as follows:
@@
While this assumption is made for the convenience of our brief introduction, it will be shown in our main theorems that this assumption is actually unnecessary; Similar arguments can be found in [3, 33].
@@ II-A Prior work
    Earlier initialization approaches typically involve two steps:
@@
    Generally, the expectation of $\mathbf{Z}$ takes the following form:
@@
    where $\alpha,\beta$ are both numerical constants, determined by the specific definition of $Z_{j}$. For instance, [27, 34] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}|\mathbf{a}_{ij}|^{2}$ with the corresponding $\alpha=1,\beta=2$, whereas [17] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}|\mathbf{a}_{ij}|$ with the corresponding $\alpha=\frac{2}{\pi\|\mathbf{x}\|},\beta\approx\frac{1}{6\|\mathbf{x}\|}$. It can be seen that there exists an apparent gap between $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})}$ and $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})^{c}}$. Specifically,
    $\min_{j\in\operatorname*{supp}(\mathbf{x})}\mathbb{E}[Z_{j}]-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}\mathbb{E}[Z_{j}]=$ $~\beta|{x_{(k)}}|^{2}$
    $\sim$ $~\frac{\|\mathbf{x}\|^{2}}{k}.$ (14)
@@
    which means one can obtain the estimate of support set by sorting the elements of $\{Z_{j}\}_{j=1}^{n}$ and selecting the indices corresponding to the $k$ largest values. This guarantees that the support estimate contains all the true candidates. Theoretical analysis shows that deriving an appropriate concentration from the gap $\frac{\|\mathbf{x}\|^{2}}{k}$ (see (14)) requires the sample complexity being [^1] When $Z_{j}$ is statistically heavy-tailed (e.g., involving the fourth power of Gaussian), some extra term containing $\log^{3}n$ (e.g., see (7)) will appear in the complexity. This is because one needs to use a truncation argument to analyze the tail bound. For simplicity, we will ignore such term throughout this subsection.
@@
    Step 2: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    The leading eigenvector of $\mathbb{E}[\mathbf{Z}]$ is $\mathbf{x}$ and $\mathbb{E}[\|\mathbf{z}\|^{2}]=\|\mathbf{x}\|^{2}$. Therefore, we shall obtain a good estimate of $\mathbf{x}$. Theoretical analysis shows that this step requires
@@
    Recently, a novel initialization paradigm was developed in [29, 3, 33], which contains four steps. Instead of directly estimating the support in one step, an index $j_{\text{max}}$ is first selected corresponding to the largest value among $Z_{j}$’s.
@@
    Establishing some concentration inequalities on this random variable yields a result similar to (15):
@@
    To identify the true support set based on (21), it requires
@@
    Step 4: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    This step is the same as Step 2 of the previous method.
@@ II-B Generalized Exponential Spectral Pursuit
As has been clarified, the enlarged gap $\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}$ leads to the improvement in complexity. Intuitively, the larger the gap is, the smaller the sample complexity will be. Note that the term $|{x_{(1)}}|$ in this gap arises from the fact that we simply select one maximal index in Step 1. If more significant indices are chosen in Step 1, the final gap will indeed be enlarged, thereby decreasing the complexity in Step 3333In fact, the complexity in Step 1 will increase while that in Step 3 will decrease as more significant indices are chosen. This means that there exists a trade-off in the number of the indices that should be chosen in Step 1.. Then the remaining issue is to design Step 2, where the set $S^{0}$ from Step 1 may contain more than one support candidate. Following the previous method, a straightforward idea is to construct the vector $\mathbf{e}^{0}$ such that it takes $1$ on $S^{0}$ and 0 for the rest. A more sophisticated design, however, is to compute the leading eigenvector of $\mathbf{Z}_{S^{0}}$ as $\mathbf{e}^{0}$. This also matches the previous method since $\mathbf{e}_{j_{\text{max}}}$ is exactly the leading eigenvector of $\mathbf{Z}_{\{j_{\text{max}}\}}$.
So far, we have completed the introduction of the proposed modification. The specific algorithm is summarized in Algorithm 1. Now we will further clarify the effectiveness intuitively. The more detailed sample complexity and numerical conclusions are formally presented in Section III.
Following [33], we utilize the spectrum:
@@
Since the leading eigenvector of $\mathbb{E}[\mathbf{Z}_{S^{0}}]$ is $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$, we can approximate $\mathbf{e}_{0}$ as $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$. Then, Step 2 ensures that
@@
This means that in Step 3, we can distinguish the true support entries. Then Step 4 can ensure that $\mathbf{z}$ is a good initialization of the target signal $\mathbf{x}$.
Algorithm 1  Generalized Exponential Spectral Pursuit
Input: $\mathbf{A}$, $\mathbf{y}$, the sparsity $k$, the parameter $p$.
@@
Step 2: Set $\mathbf{e}^{0}$ as the unit leading eigenvector of $\mathbf{Z}_{S^{0}}$.
@@
Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.